Mr. Arroyo bought a piece of property for P100,000 down payment and 10 deferred semiannual payments of P8,000 starting 3 years from now. If the interest rate is 12% compounded semiannually, what is the present worth of the property?





1
Expert's answer
2022-03-20T19:06:15-0400

"P_1=\frac{A[(1+i)]^{n}-1}{(1+i)^{n}i}=\frac{8000[(1+0.06)]^{10}-1}{(1+0.06)^{10}(0.06)}=58,880.69\\P_1=P_2(1+i)^n\\58.860.69=P_2(1+0.06)^5=43,999.078\\Total amount=100,000+P_2\\=100,000+43,999.078\\=143,999.078"